Neptune Company produces toys and other items for use in beach and resort areas. A small, inflatable toy has come onto the market that the company is anxious to produce and sell. The new toy will sell for $3.40 per unit. Enough capacity exists in the company’s plant to produce 30,500 units of the toy each month. Variable expenses to manufacture and sell one unit would be $2.14, and fixed expenses associated with the toy would total $57,145 per month.

The company's Marketing Department predicts that demand for the new toy will exceed the 30,500 units that the company is able to produce. Additional manufacturing space can be rented from another company at a fixed expense of $2,857 per month. Variable expenses in the rented facility would total $2.38 per unit, due to somewhat less efficient operations than in the main plant.

Required:

1. What is the monthly break-even point for the new toy in unit sales and dollar sales.

2. How many units must be sold each month to attain a target profit of $13,056 per month?

3. If the sales manager receives a bonus of 25 cents for each unit sold in excess of the break-even point, how many units must be sold each month to attain a target profit that equals a 25% return on the monthly investment in fixed expense-29
Answer #1

Contribution per unit from company = $3.4 $ 2.14 = $ 1.26

Contribution per unit from additional manufacturing space= $3.40 - $ 2.38 = $ 1.02

1) monthly break - even point = Fixed expenses / contribution per unit

= 57145/ 1.26

= 45,353 units

Since the breakeven point in greater than the company 's capacity, additional manufacturing space should be rented.

Here, Then breakeven point = ( total fixed expenses- contribution from units from primary capacity of company) / (contribution per unit in new additional manufacturing space)

= [(57,145 + 2,857) - ( 30500 units * 1.26 ) ] / 1.02

= (60,002 - 38,430) / 1.02

= 21,572 / 1.02

= 21,149 units

Total breakeven point = 30,500 units + 21,149 units

= 51,649 units

Total breakeven point (dollar) = 51649 units * $ 3.40

= $ 1,75,606.60

Hence, monthly breakeven point (unit sales) = 51,649 units

monthly breakeven point ( dollar) = $ 1,75,606.60

(2)Amoest (6) Pasticales Target Profit Total fixed expesse 13,056 60,00a. 73058 Total Lesso Contributors fooso ceoits from prima

Hence, number of units to be sold each month to attain a target profit of $ 13,056 per month = 33,949 units

(3)

Particalars Account ] Target Profit Required. 60,002 x 250] 15.000.50 Fixed ecopesse. (57,145 + 2,851) Goooa 75.002.50 Less:

Hence, number of units to be sold each month to attain a target profit= 19481 units